please help me solve this word problem which has to do with using formulas in geometry. find the length of the sides of a square whose area and perimeter are the same nonzero number
Answer by mananth(16946)


let the length of the side be L
Area of square = L*L = L^2
perimeter of square = 4L
...
they are same
L^2=4L
L^2/L =4
L=4